OpenStudy (mhchen):

Rule of Means: Proper Notation shown below:

OpenStudy (mhchen):

\[\mu_{x-y} = \mu_{y-x}?\] I came across this question. I'm pretty sure it's not but I just want to make sure

OpenStudy (reemii):

What is the context?

OpenStudy (mhchen):

Context doesn't matter, it's just the general rule.

OpenStudy (reemii):

I don't understand.. are you talking about probability distributions (gaussian?) ? You should define what is \(\mu_a\) .. for some \(a\)..

OpenStudy (amorfide):

I suck at stats but does this help you at all http://www.kaspercpa.com/statisticalreview.htm

OpenStudy (mhchen):

I checked that website actually. It stated that E(X+Y) = E(X) + E(Y), didn't say anything about E(X-Y) though so I guess they thought it would be E(X) - E(Y) Okay sorry about the context. It's for a simple probability distribution.

OpenStudy (amorfide):

honestly I have no idea about stats but E(X-Y)=E(X+(-Y) let Z=-Y E(X+Z)=E(X)+E(Z) so E(X+Z)=E(X)+E(-Y) and E(aX)=aE(X) so E(-Y)=-1E(Y) so E(X-Y)=E(X)-E(Y) am I wrong??

OpenStudy (amorfide):

@mhchen

OpenStudy (reemii):

@amorfide Your development is correct. Out of the blue, I would say: rule of means is wrong the rule with the sigma's is correct. (variance does not suffer from a negative sign)

OpenStudy (mhchen):

Yeah I got it. You proved E(X-Y) = E(X) - (Y) So based on that E(X-Y) does NOT equal to E(Y-X)

OpenStudy (amorfide):

E(X-Y)=E(-(Y-X)) let Z=Y-X E(-Z)=-E(Z) hence E(X-Y)=-E(Y-X) I think

OpenStudy (mhchen):

yep.